OpenStudy (anonymous):

Find the roots of x^3 – x^2 + 2 = 0

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

um not really

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Try +/- 1 and +/- 2. Always check ro Rational Roots. If you manage to find one, the world is a better place!

OpenStudy (anonymous):

so x = –1 + 2i, x = –1, x = – 2i, x = 1

OpenStudy (tkhunny):

Really? Four roots from a Cubic Equation? Something fishy about that.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

1. Find the roots of x3 – x2 + 2 = 0. A. x = 1 + i, x = 1 – i, x = –1 B. x = –1 + 2i, x = –1, x = – 2i, x = 1 C. x = –1 + i, x = –1 – i, x = –1 D. x = 1 + 2i, x = 1 – 2i, x = 1

OpenStudy (anonymous):

see

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Okay, then discard B immedaitely. It absolutely cannto be correct. Now, quit guessing. Did you yet find a factoring? A root? Prove it.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

it means x-1=0

OpenStudy (anonymous):

now plug x=-1

OpenStudy (anonymous):

if it gives u ...zero ...then it is ur root

OpenStudy (anonymous):

i mean first root

OpenStudy (anonymous):

...... aye, eh, thank but im thinking its C

OpenStudy (tkhunny):

Thinking? That's just not good enough. Prove it!! x^3 – x^2 + 2 = 0 Rational Roots can be ONLY +/- 1, +/- 2 A quick inspection says there MUST be a Negative Real Root. Maybe it's rational! Let's try x = -1 (-1)^3 - (-1)^2 + 2 = -1 -1 + 2 = 0 -- That's it!!!! x = -1 is a root. (x+1) is a factor. Now we reduce the problem (x+1)(x^2 - 2x + 2) = 0 That last part doesn't factor. Use the Quadratic Formula to find the other TWO roots. Quadratic ==> Two Cubic ==> Three Quartic ==> Four Quintic ==> Five

OpenStudy (anonymous):

guyes theres more than one answer

OpenStudy (anonymous):

A. x = 1 + i, x = 1 – i, x = –1 B. x = –1 + 2i, x = –1, x = – 2i, x = 1 C. x = –1 + i, x = –1 – i, x = –1 D. x = 1 + 2i, x = 1 – 2i, x = 1 see????

OpenStudy (nincompoop):

x^3-x^2+2 = 0 Factor the left hand side. The left hand side factors into a product with two terms: (x+1) (x^2-2 x+2) = 0 Solve each term in the product separately. Split into two equations: x+1 = 0 or x^2-2 x+2 = 0 Look at the first equation: Solve for x. Subtract 1 from both sides: x = -1 or x^2-2 x+2 = 0 Look at the second equation: Solve the quadratic equation by completing the square. Subtract 2 from both sides: x = -1 or x^2-2 x = -2 Take one half of the coefficient of x and square it, then add it to both sides. Add 1 to both sides: x = -1 or x^2-2 x+1 = -1 Factor the left hand side. Write the left hand side as a square: x = -1 or (x-1)^2 = -1 Eliminate the exponent on the left hand side. Take the square root of both sides: x = -1 or x-1 = i or x-1 = -i Look at the second equation: Solve for x. Add 1 to both sides: x = -1 or x = 1+i or x-1 = -i Look at the third equation: Solve for x. Add 1 to both sides: Answer: x = -1 or x = 1+i or x = 1-i

OpenStudy (anonymous):

x=0, 1 and 1/2+root7/2 i

OpenStudy (anonymous):

thats not an optoin!

OpenStudy (tkhunny):

B. No good. There are four. There can be only 3. D. No good. Does not include x = -1 @selenamalter You have shown NO work of your own. You cannot solve ANY problem with that kind of effort. Don't GUESS at the answer. Prove it!

OpenStudy (anonymous):

(x+1) (x^2-2x+2)=0 solve it

OpenStudy (anonymous):

x=-1 ...for second part...use quardatic eqaution

OpenStudy (anonymous):

x=-1, x= 2+i/2 x=2-i/2

OpenStudy (anonymous):

yeah u can get that

OpenStudy (anonymous):

do u rem the formula for quardatic equation

OpenStudy (anonymous):

it comes like 2+0r - square root of 4 i over 2

OpenStudy (anonymous):

do it u will get that answer

OpenStudy (tkhunny):

Seriously, @basa, quit typing! Serously, @selenamalter, show your work. What is your plan for a solution?

OpenStudy (anonymous):

He has none. He's just guessing.

OpenStudy (anonymous):

But I think that was obvious.

OpenStudy (anonymous):

...... ya confusing meh... oh um

OpenStudy (anonymous):

ugh! forget it i quit

OpenStudy (tkhunny):

This will almost always be the result when you do no work of your own.

OpenStudy (anonymous):

STHAP IT I AM eh, mabe i eh i dont undertand hmmm?

OpenStudy (anonymous):

yeh only possible answes are x=-1,x=1+i,x=1-i

OpenStudy (anonymous):

you better plug those values in the equations and check them out..it should give u zero..n it does

OpenStudy (tkhunny):

@ selenamalter I believe that you do have some clue. You have shown it on other problems in other threads. You just didn't show any of it in this thread. Not anything. Go back and look. There is nothing. That will never be good enough.

OpenStudy (anonymous):

yep total of real and complex solutions as following: Real: x=-1 Imag: x=1-1i,1+1i

OpenStudy (anonymous):

x=-1.....it makes that equation zero so it is one of the roots of that eqations..do same for other two

OpenStudy (anonymous):

x^3+x^2-2x^2-2x+2x+2=0 imples x^2(x+1)-2x(x+1)+2(x+1)=0 implies (x+1)(x^2-2x+1)=0